OpenStudy (anonymous):

Determine the Laplace transform\[\large e^{-t} \space t \space \sin2t\]\[f(t)=t \space \sin2t\]\[L[e^{-t}f(t)]=L[f(t)](s+1)\]My Answer\[\frac{ 4s(s+1) }{ (s^2+4)^2 }\]Book Answer\[\frac{ 4(s+1) }{ [(s+1)^2+4]^2 }\]What am I missing?

OpenStudy (anonymous):

@zepdrix You feel like looking at another one?

zepdrix (zepdrix):

Hmmm I dunno, I gotta get caught up on my Diff EQ homework apparently XD lol

OpenStudy (anonymous):

I have a Laplace test tomorrow and I'm just going back thru all the study problems. I feel like I did this right. I don't know. What year are you in school @zepdrix

zepdrix (zepdrix):

Mmmmm i dunno, almost done with 2 years i guess :O not too far in yet.

OpenStudy (anonymous):

your major?

zepdrix (zepdrix):

Pshhhh i dunno, I'll figure that out next semester.. trying to put it off as long as i can, until i absolutely have to decide XD I love love love math.. i can't seem to narrow it down any further than that :D

OpenStudy (anonymous):

Haha, took me 4 years and a couple of different majors before I landed on engineering. I might post more Laplace questions. stay tuned

OpenStudy (accessdenied):

Here is what I did to get the book answer: http://mathbin.net/114299

OpenStudy (anonymous):

@AccessDenied Thank you for the solution. I'll study that for future problems :)

OpenStudy (anonymous):

simplify \[(\sqrt x+\sqrt 3)(\sqrt x+\sqrt 27)\]\[(\sqrt x \times \sqrt x)+(\sqrt x \times \sqrt 27)+(\sqrt 3 \times \sqrt x)+(\sqrt 3 \times \sqrt 27)\]\[x+\sqrt {27x}+\sqrt {3x}+[\sqrt {3 \times 27}=\sqrt {81}=9]\]@lala2

OpenStudy (anonymous):

thanks so much @ChmE!

OpenStudy (anonymous):

yup